OpenStudy (anonymous):

Find the sums and show work 1+2+3+ ... +3218+3219+3220

OpenStudy (anonymous):

you just add it up and it equals to 9663

jhonyy9 (jhonyy9):

mke the same like in case of 500 1+3219=3220 2+3218=3220 3+3217=3220 ...

jhonyy9 (jhonyy9):

make this till the end so after you need calcul how many from 3220 there are multiply this number with 3220 and will get the result

OpenStudy (anonymous):

3220(3221)/2 =5185810